OpenStudy (anonymous):

I will give medal! 12(20 − 17) − 3 ∙ 6 i need someone to help me understand.

OpenStudy (anonymous):

Do you know PEMDAS?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Please excuse my dear aunt sally?

OpenStudy (anonymous):

ooh ya

OpenStudy (anonymous):

P Paratheses E Exponets M Multiplication D Division A Addition S Subtraction All multiplication and division from left to right All addition and subtraction from left to right also.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know how to solve the problem?;o

OpenStudy (anonymous):

mmm

OpenStudy (anonymous):

im not getting it still

OpenStudy (anonymous):

12(20 − 17) − 3 ∙ 6 P Paratheses E Exponets M Multiplication D Division A Addition S Subtraction Step 1 Paratheses - (20 - 17) = ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

12(20 − 17) − 3 ∙ 6 P Paratheses E Exponets M Multiplication D Division A Addition S Subtraction Step 1 Paratheses - (20 - 17) = 3 Between 12(20 always add a multiplication sigh like " 12 x (20

OpenStudy (anonymous):

12(20 − 17) − 3 ∙ 6 P Paratheses E Exponets M Multiplication D Division A Addition S Subtraction Step 1 Paratheses - (20 - 17) = ?

OpenStudy (anonymous):

12(20 − 17) − 3 ∙ 6 P Paratheses E Exponets M Multiplication D Division A Addition S Subtraction Step 1 Paratheses - (20 - 17) = ?

OpenStudy (anonymous):

ooh i remember so you always start with the ( ) i remember this in 6 grade when i was in school

OpenStudy (anonymous):

Its okay im in 7th grade xD

OpenStudy (anonymous):

lol you probably would be shocked but im in 9 and trying to get in my right grade which is 10th

OpenStudy (anonymous):

so its 18

OpenStudy (anonymous):

You got it!!! and its okay xD School sucks

OpenStudy (anonymous):

ya alot thanks :)

OpenStudy (anonymous):

Tag me if you need anymore help @Pizzalover123

OpenStudy (anonymous):

Your welcome xD Anytime

OpenStudy (anonymous):

oki